OpenStudy (anonymous):

The absorbances of a series of solutions containing Cr3+ were measured at a fixed wavelength and a path length of 1.1 cm. The following data were recorded: 4.0 mL of 0.100 M Cr3+ solution; 1.0 mL of water; Absorbance = 0.807; (three more solutions go like this in same order as listed previously: 3.0, 2.0, 0.608; 2.0, 3.0, 0.394; and 1.0,4.0, 0.193. Calculate the concentration of Cr3+. please explain in detail, i'm a beginner. also: someone said to plot this data to get the slope which is e in the beer-lambert law: A=eCl. but what do i put in the x and y axis? im very confused.

OpenStudy (cuanchi):

I think the problem is missing the last information about the absorbance of the Cr3+ solution of which you would like to know the concentration. With the data from your table (your standard curve) you can calculate the molar absorbtivity of the Cr+3. that is the e in the Beer Lambert law (as you said is the slope of the line in the graph) or can be calculated mathematically as the average of the quotient between the A/C (absorbance divided concentration) Off course you need to calculate the concentration of your standards 4mL O.1M + 1mL water = ??? M M1 x C1 = M2 x C2 M2 = V1 x C1 / V2 = 4 mL x 0.1 M / 5 mL = 0.08M This is the molar concentration of your first standard point then you can do the same for the other 3 standards then divide the absorbance by the concentration 0.807 / 0.08 = 10.08 (this is "e") for the first point. repeat the same for the other 3 concentration and you will get numbers around 10.08 to 9.65. Calculate the average of these 4 values and that is your average "e" Now is when you need the data that is not given in the problem that is the absorbance of the unknown solution. If you have it then replace that in the Beer Lambert equation and calculate the concentration whit the value of average "e" that you calculate from your standards.